Hello. May I have some help solving this equation?
Find all real solutions of the following equation. (If there are extra answer boxes, enter NONE.)
x^4 - 13x2 + 30 = 0

Answer by checkley71(8403)

X^4-13X^2+30=0
(X^2-10)(X^2-3)=0
X^2-10=0
X=SQRT10
X=3.16 ANSWER
X^2-3=0
X^2=3
X=1.73 ANSWER.